# NeCo[Uncaptioned image]: Improving DINOv2’s spatial representations in 19 GPU hours with Patch Neighbor Consistency

### v1

@@ 4.2 Comparison to State-of-the-Art
Figure 2: In-context scene understanding benchmark. Dense nearest neighbor retrieval performance is reported across various training data proportions on two scene-centric datasets, Pascal VOC and ADE20k. The retrieved cluster maps are compared with the ground truth using Hungarian matching [57], and their mIoU score is reported. For all models, ViT-S16 is used except for DINOv2R and NeCo, where it is ViT-S14. For full tables, refer to Appendix B.3
@@
In this section, we first compare the quality of frozen features learned through NeCo  with state-of-the-art methods in in-context learning via nearest neighbor retrieval and unsupervised semantic segmentation tasks. Next, we demonstrate that NeCo’s versatility by applying it to five different pretraining models and show it improves their dense features consistently. Finally, we evaluate the transferability of our learned dense representations to other datasets by using linear head semantic segmentation and end-to-end fine-tuning with Segmentor [48].
Visual In-Context Learning Evaluation.
@@
We compare our approach to a recently proposed benchmark [10] that evaluates in-context reasoning in vision models. Unlike traditional linear segmentation methods, this evaluation does not require fine-tuning or end-to-end training. Instead, it creates validation segmentation maps by matching patch-level, feature-wise nearest neighbor similarities between validation images (queries) and training samples (keys). This method, inspired by NLP strategies, tests how well models learn tasks from a few examples. The results are presented in Figure 2. As shown, NeCo outperforms prior state-of-the-art methods such as CrIBo and DINOv2R by 4% to 13% on Pascal and ADE20k across different fractions. The performance gap between NeCo and others increases, particularly in the data-efficiency regime. This improvement is due to our method’s explicit enforcement of patch-level nearest neighbor consistency, resulting in higher-quality patch-level representations that remain effective even with fewer images. In contrast, other methods that promote image-level [10] or object-level [9] consistency force consistency between a pooled vector of patches, potentially leading to inadequate semantic patch-level representations, particularly in smaller datasets. Our method’s ability to perform well with few images brings vision models one step closer to in-context learning style generalist reasoning. For complete tables and details on scaling the results to larger models, please refer to Appendix B.3. For visualizations please refer to Appendix C.
@@
Table 1: Frozen clustering-based evaluations. (a) We evaluate the models by running $K$-means with various clustering granularities $K$ on the spatial features on two datasets. The resulting cluster maps are matched to the ground-truth by Hungarian matching, and the intersection is reported in mIoU. (b) Following previous works [1, 2], we post-process the resulting maps and report unsupervised semantic segmentation on Pascal VOC. Both tables use ViT-S with the patch size of 16, except for DINOv2 and NeCo, where it is 14.
@@
DINO [13] 4.3 13.9 17.3 5.4 18.8 19.2
iBOT [14] 4.4 23.8 31.1 7.6 26.6 28.0
CrOC [4] 3.4 16.4 20.0 4.9 14.7 18.1
TimeT [2] 12.2 43.6 46.2 17.5 42.7 44.6
DINOv2R [15] 12.2 46.7 49.5 12.3 38.9 41.2
CrIBo [9] 18.3 51.3 54.5 14.5 46.0 48.3
@@
MaskConstrast [19] 35.1
DINOv2R [15] 35.1
DeepSpectral [58] 37.2
DINOSAUR [37] 37.2
Leopart [1] 41.7
COMUS [44] 50.0
@@
iBOT [14] 4.4 31.1 66.1 15.4↑11.0 51.2↑20.1 68.6↑2.5 7.6 28.0 58.9 20.4↑12.8 52.8↑24.8 67.7↑8.8
DINO [13] 4.3 17.3 50.2 14.5↑10.2 47.9↑30.6 61.3↑11.1 5.4 19.2 43.9 16.9↑11.5 50.0↑30.8 62.4↑18.5
TimeT [2] 12.2 46.2 66.3 17.9↑5.7 52.1↑5.9 68.5↑2.2 18.4 44.6 58.2 20.6↑2.2 54.3↑9.7 64.8↑6.6
Leopart [1] 15.4 51.2 66.5 21.0↑5.6 55.3↑4.1 68.3↑1.8 14.8 53.2 63.0 18.8↑4.0 53.9↑0.7 65.4↑2.4
CrIBo [9] 18.3 54.5 71.6 21.7↑3.4 59.6↑5.1 72.1↑0.5 14.5 48.3 64.3 21.1↑6.6 54.0↑5.7 68.0↑3.7
Frozen Clustering-based Evaluations.
Next, we evaluate the representation quality of our learned dense features across different objects in each dataset. Ideally, we expect that all patch features belonging to the same object, when clustered, will be assigned to the same cluster. If the learned representations are more fine-grained, such as learning object parts instead of whole objects (e.g., hands or faces instead of a person), they should consistently cover the same part across the entire dataset. To measure this, we extract dense features from all images and apply $K$-means clustering with various $K$ values to create cluster maps for each image. These cluster maps are then matched with the ground truth using Hungarian matching [49], and their mIoU is reported. For the first scenario, $K$ matches the number of ground truth objects. Additionally, to account for the second scenario, we also report performance in overclustering setups.
The results in 1(a) show that NeCo passes state-of-the-art by CrIBo by 14.5% on average across various datasets and metrics. Note that this gain is not due to the DINOv2R initialization, as it performs 4% lower than CrIBo on average. In 1(b), we report clustering performance when $K$ matches the number of ground truth objects, with clustering applied solely to foreground patches extracted by methods used in [1, 2]. We outperform other methods by at least 5.1% without relying on self-training, which requires training a separate segmentation head, as used in COMUS.
Linear Semantic Segmentation Evaluation.
In this experiment, we keep the pretrained backbone frozen and train a linear layer on top of the spatial features to solve a supervised semantic segmentation task. Bilinear interpolation is used to match the spatial feature resolution to the image size, enabling the application of pixel-wise cross-entropy loss. This setup provides a better evaluation of the pretrained models compared to end-to-end finetuning, where all learned parameters are overwritten. The results, reported in Table 2, show that NeCo surpasses CrIBo on all datasets by at least 10% and outperforms DINOv2R by 5% to 7%. These significant improvements demonstrate that patches representing the same object or object part have higher similarities in feature space compared to other methods, as a simple linear layer can utilize these features for strong semantic segmentation.
Compatibility with Differently Pretrained Backbones.
As shown in Table 3, our method is generalizable across various self-supervised learning initialization, improving them by roughly 4% to 30% across different metrics and datasets. Surprisingly, NeCo even enhances the performance of methods specifically designed for dense tasks, such as CrIBo, TimeT, and Leopart. For instance, CrIBo demonstrates a performance increase of approximately 5% in overclustering evaluations, which measures how fine-grained and semantic the representations learned during pretraining are. This indicates that NeCo applied to CrIBo can extract more discriminative features, leading to improved transfer performance, shown by 0.5% and 3.7% better linear classification performance on Pascal VOC and COCO-Things.
End-to-End Full-Finetuning Evaluation.
@@
One advantage of self-supervised pretraining is the ability to transfer learned general semantic features to specialized downstream tasks, improving performance in an end-to-end finetuning setup. We evaluate this capability of NeCo  by adding a transformer-based decoder from Segmenter [48] on top of the feature extractor and finetuning the entire network for semantic segmentation. The backbone’s spatial features are fed into a transformer decoder along with $K$ learnable class tokens. These class and spatial tokens are projected onto each other to obtain patch-level predictions, which are then upsampled to match the input image size, enforcing pixel-wise cross-entropy loss. We report the mIoU scores achieved on Pascal VOC, Pascal Context, COCO-Stuff, and ADE20k in Table 4. Despite all parameters being adapted, the results show that NeCo  learns superior features, leading to better performance in downstream tasks, outperforming CrIBo by approximately 4%. Notably, while DINOv2R has demonstrated strong transfer results in various tasks, including semantic segmentation, due to being trained on a massive dataset of 142M images and using a combination of dense and classification losses, NeCo surpasses even this. By training only 19 GPU-hours on COCO, which is a fraction of the original compute, we obtain consistent gains, setting a new state-of-the-art.
@@
Table 4: Evaluation of Full Fine-Tuning with Segmenter. Various backbones pre-trained with different self-supervised learning methods are fine-tuned using Segmenter [48]. The table shows the mIoU scores obtained on validation sets across 4 different datasets.
@@
Table 5: Ablations of the key parameters of our method. We evaluate the models by training a linear layer on top of the frozen representations (Lin.) or using the in-context (IC) evaluation of [10] using the validation images for PascalVOC12 and ADE20k.
@@
Method Lin. IC Lin. IC
Odd-even 78.9 62.2 37.0 21.6
Bitonic 78.9 62.0 37.3 21.7
4.3 Ablation Studies
Here, we examine the essential parameters of our method by training NeCo on Pascal VOC12 and ADE20k. We assess its ability to perform linear segmentation and in-context scene understanding using the frozen representations learned with each set of parameters. For in-context scene understanding evaluations, we use $\frac{1}{128}$ fraction of the training data and reduce the spatial dimension to $448^{2}$. The number of training epochs for linear segmentation evaluations is set to 20 epochs. For more ablations refer to Appendix B.
Patch Selection Approach.
We demonstrate the effect of selecting patches from the foreground, background, or both in 5(a). Foreground patches are selected using the attention map averaged across heads. Our results indicate that selecting patches from the foreground gives 1% better results compared to the background selection in 3 out of 4 metrics. However, the performance peaks when we select patches from both locations. This improvement can be attributed to the use of scene-centric images for training, where the background often contains meaningful objects that contribute to enhanced performance.
Utilizing a Teacher.
We ablate the role of teacher-student architecture in 5(b). As shown, employing a teacher network updated by exponential moving average can significantly improve the performance across all the metrics by 8% to 20%. This is consistent with the previous works [13, 29], which reported a more stable training process when the teacher-student architecture is employed.
Nearest Neighbor Selection Approach.
We evaluate the influence of picking nearest neighbors from the same image (intra) or different batch images (inter) in 5(c). The results indicate that selecting patches across images consistently boosts performance by roughly 0.4% to 1% across different metrics. The higher diversity of patches involved in the latter approach likely accounts for this improvement.
Training Dataset.
5(d) presents the impact of the training dataset based on the ImageNet-100, Pascal, and COCO datasets. ImageNet-100 comprises relatively simple images with few objects, whereas Pascal and COCO feature more complex scenes with multiple objects. Our method shows consistent improvements when trained on multi-object datasets, achieving a performance increase of 2% to 7% on COCO compared to ImageNet-100. This improvement is due to the greater quantity and diversity of objects per batch in multi-object scenes, which provide stronger learning signals by requiring discrimination against a higher number of objects. Notably, the additional performance boost we observe from finetuning DINOv2R on Pascal—despite it already being trained on this dataset [15]—further underscores the efficacy of our proposed loss function.
Sorting Algorithm.
We ablate the effect of changing the sorting algorithm in 5(e). Our method demonstrates robust performance across different sorting approaches, achieving the best results with Bitonic sorting, which yields slightly better performance on average. We present the ablation study on the sorting steepness parameter in the Appendix B.4, as our method is robust to variations in this parameter.
5 Discussion
Broader impact.
By leveraging pretrained models, our method has implications for reducing the carbon footprint of model training. Similar to how once an LLMs is trained, it can be re-used and tuned with very little compute, we demonstrate that this is possible for visual foundation models. Our finetuning requires only 19 GPU-h, yet yields vastly better models for dense image understanding. We believe this work spurs further research into post-pretraining methods.
Conclusion.
In this work, we propose Patch Nearest Neighbor Consistency as a new method for dense post-pretraining of self-supervised backbones. By applying our method to the many backbones including the DINOv2-registers model, we improve upon these models by a large margin for frozen clustering, semantic segmentation and full finetuning, setting several new state-of-the-art performances.

### v2

@@ 4.2 Comparison to State-of-the-Art
Figure 2: In-context scene understanding benchmark. Dense nearest neighbor retrieval performance is reported across various training data proportions on two scene-centric datasets, Pascal VOC and ADE20k. The retrieved cluster maps are compared with the ground truth using Hungarian matching (Kuang et al., 2021), and their mIoU score is reported. For all models, ViT-S16 is used except for DINOv2R and NeCo, where it is ViT-S14. For full tables, refer to Appendix B.3
@@
In this section, we first compare the quality of frozen features learned through NeCo  with state-of-the-art methods in in-context learning via nearest neighbor retrieval and unsupervised semantic segmentation tasks. Next, we demonstrate NeCo’s versatility by applying it to five different pretraining models and show it improves their dense features consistently. We then evaluate the transferability of our learned dense representations to other datasets by using linear head semantic segmentation and end-to-end fine-tuning with Segmentor (Strudel et al., 2021). Finally, we show that NeCo  improves the 3D understanding of different vision models, as proposed in  El Banani et al. (2024), using the multiview consistency experiment.
@@
Visual In-Context Learning Evaluation. We compare our approach to a recently proposed benchmark (Balazevic et al., 2023) that evaluates in-context reasoning in vision models. Unlike traditional linear segmentation methods, this evaluation does not require fine-tuning or end-to-end training. Instead, it creates validation segmentation maps by matching patch-level, feature-wise nearest neighbor similarities between validation images (queries) and training samples (keys). This method, inspired by NLP strategies, tests how well models learn tasks from a few examples. The results are presented in Figure 2. As shown, NeCo outperforms prior state-of-the-art methods such as CrIBo and DINOv2R by 4% to 13% on Pascal and ADE20k across different fractions. The performance gap between NeCo and others increases, particularly in the data-efficiency regime. This improvement is due to our method’s explicit enforcement of patch-level nearest neighbor consistency, resulting in higher-quality patch-level representations that remain effective even with fewer images. In contrast, other methods that promote image-level (Balazevic et al., 2023) or object-level (Lebailly et al., 2023) consistency force consistency between a pooled vector of patches, potentially leading to inadequate semantic patch-level representations, particularly in smaller datasets. Our method’s ability to perform well with few images brings vision models one step closer to in-context learning style generalist reasoning. For complete tables and details on scaling the results to larger models, please refer to Appendix B.3. For visualizations please refer to Appendix C.
@@
Table 1: Frozen clustering-based evaluations. (a) We evaluate the models by running $K$-means with various clustering granularities $K$ on the spatial features on two datasets. The resulting cluster maps are matched to the ground-truth by Hungarian matching, and the intersection is reported in mIoU. (b) Following previous works (Ziegler & Asano, 2022; Salehi et al., 2023), we post-process the resulting maps and report unsupervised semantic segmentation on Pascal VOC. Both tables use ViT-S with the patch size of 16, except for DINOv2R and NeCo, where it is 14.
@@
DINO 4.3 13.9 17.3 5.4 18.8 19.2
iBOT 4.4 23.8 31.1 7.6 26.6 28.0
CrOC 3.4 16.4 20.0 4.9 14.7 18.1
TimeT 12.2 43.6 46.2 17.5 42.7 44.6
DINOv2R 12.2 46.7 49.5 12.3 38.9 41.2
CrIBo 18.3 51.3 54.5 14.5 46.0 48.3
@@
MaskConstrast 35.1
DINOv2R 35.1
DeepSpectral 37.2
DINOSAUR 37.2
Leopart 41.7
COMUS 50.0
@@
Frozen Clustering-based Evaluations.
Next, we evaluate the representation quality of our learned dense features across different objects in each dataset. Ideally, we expect that all patch features belonging to the same object, when clustered, will be assigned to the same cluster. If the learned representations are more fine-grained, such as learning object parts instead of whole objects (e.g., hands or faces instead of a person), they should consistently cover the same part across the entire dataset. To measure this, we extract dense features from all images and apply $K$-means clustering with various $K$ values to create cluster maps for each image. These cluster maps are then matched with the ground truth using Hungarian matching (Kuhn, 1955), and their mIoU is reported. For the first scenario, $K$ matches the number of ground truth objects. Additionally, to account for the second scenario, we also report performance in overclustering setups.
The results in 1(a) show that NeCo passes state-of-the-art by CrIBo by 14.5% on average across various datasets and metrics. Note that this gain is not due to the DINOv2R initialization, as it performs 4% lower than CrIBo on average. In 1(b), we report clustering performance when $K$ matches the number of ground truth objects, with clustering applied solely to foreground patches extracted by methods used in Ziegler & Asano (2022); Salehi et al. (2023). We outperform other methods by at least 5.1% without relying on self-training, which requires training a separate segmentation head, as used in COMUS. For visualizations please refer to Appendix C.
Linear Semantic Segmentation Evaluation. In this experiment, we keep the pretrained backbone frozen and train a linear layer on top of the spatial features to solve a supervised semantic segmentation task. Bilinear interpolation is used to match the spatial feature resolution to the image size, enabling the application of pixel-wise cross-entropy loss. This setup provides a better evaluation of the pretrained models compared to end-to-end finetuning, where all learned parameters are overwritten. The results, reported in Table 2, show that NeCo surpasses CrIBo on all datasets by at least 10% and outperforms DINOv2R by 5% to 7%. These significant improvements demonstrate that patches representing the same object or object part have higher similarities in feature space compared to other methods, as a simple linear layer can utilize these features for strong semantic segmentation.
@@
Compatibility with Differently Pretrained Backbones. As shown in Table 3, our method is generalizable across various self-supervised learning initialization, improving them by roughly 4% to 30% across different metrics and datasets. Surprisingly, NeCo even enhances the performance of methods specifically designed for dense tasks, such as CrIBo, TimeT, and Leopart. For instance, CrIBo demonstrates a performance increase of approximately 5% in overclustering evaluations, which measures how fine-grained and semantic the representations learned during pretraining are. This indicates that NeCo applied to CrIBo can extract more discriminative features, leading to improved transfer performance, shown by 0.5% and 3.7% better linear classification performance on Pascal VOC and COCO-Things.
@@
iBOT (Zhou et al., 2022) 4.4 31.1 66.1 15.4↑11.0 51.2↑20.1 68.6↑2.5 7.6 28.0 58.9 20.4↑12.8 52.8↑24.8 67.7↑8.8
DINO (Caron et al., 2021) 4.3 17.3 50.2 14.5↑10.2 47.9↑30.6 61.3↑11.1 5.4 19.2 43.9 16.9↑11.5 50.0↑30.8 62.4↑18.5
TimeT (Salehi et al., 2023) 12.2 46.2 66.3 17.9↑5.7 52.1↑5.9 68.5↑2.2 18.4 44.6 58.2 20.6↑2.2 54.3↑9.7 64.8↑6.6
Leopart (Ziegler & Asano, 2022) 15.4 51.2 66.5 21.0↑5.6 55.3↑4.1 68.3↑1.8 14.8 53.2 63.0 18.8↑4.0 53.9↑0.7 65.4↑2.4
CrIBo (Lebailly et al., 2023) 18.3 54.5 71.6 21.7↑3.4 59.6↑5.1 72.1↑0.5 14.5 48.3 64.3 21.1↑6.6 54.0↑5.7 68.0↑3.7
@@
End-to-End Full-Finetuning Evaluation. One advantage of self-supervised pretraining is the ability to transfer learned general semantic features to specialized downstream tasks, improving performance in an end-to-end finetuning setup. We evaluate this capability of NeCo  by adding a transformer-based decoder from Segmenter (Strudel et al., 2021) on top of the feature extractor and finetuning the entire network for semantic segmentation. The backbone’s spatial features are fed into a transformer decoder along with $K$ learnable class tokens. These class and spatial tokens are projected onto each other to obtain patch-level predictions, which are then upsampled to match the input image size, enforcing pixel-wise cross-entropy loss. We report the mIoU scores achieved on Pascal VOC, Pascal Context, COCO-Stuff, and ADE20k in Table 4. Despite all parameters being adapted, the results show that NeCo  learns superior features, leading to better performance in downstream tasks, outperforming CrIBo by approximately 4%. Notably, while DINOv2R has demonstrated strong transfer results in various tasks, including semantic segmentation, due to being trained on a massive dataset of 142M images and using a combination of dense and classification losses, NeCo  surpasses even this. By training only 19 GPU-hours on COCO, which is a fraction of the original compute, we obtain consistent gains, setting a new state-of-the-art.
Multiview Feature Consistency Evaluation. This evaluation assesses the 3D understanding of models through geometric correspondence estimation, aiming to measure the consistency of feature representations across multiple views of the same scene. By identifying matching points in different images without additional model training, it provides a direct evaluation of 3D feature quality (El Banani et al., 2024). The evaluation is done on a large-scale dataset called Spair-71k (Min et al., 2019). As Table 5 shows, NeCo  can boost the performance of DINO models by roughly 10% on average across different categories. This demonstrates that the proposed method extends beyond enhancing semantic segmentation and has a broader impact on vision foundation models by improving their overall spatial understanding
@@
Table 4: Evaluation of Full Fine-Tuning with Segmenter. Various backbones pre-trained with different self-supervised learning methods are fine-tuned using Segmenter (Strudel et al., 2021). The table shows the mIoU scores obtained on validation sets across 4 different datasets.
@@
Table 5: Multiview feature consistency results on SPair-71k, Recall@0.01. We use the evaluation method proposed by El Banani et al. (2024). NeCo  improves the results of DINO models by roughly 1.8% for 3D understanding, measured by multiview feature consistency.
[…]
DINO-B16 26.7 14.9 35.4 15.6 20.6 19.3 18.1 33.7 11.2 19.4 23.9 16.3 16.3 18.8 11.1 12.9 29.5 10.1 19.65
+ NeCo 26.0 17.5 36.7 16.9 22.0 23.5 18.0 33.5 10.6 21.0 21.6 20.2 20.6 19.3 13.7 13.9 36.3 12.5 21.32
DINOV2R-B14 31.2 34.4 56.7 23.2 26.1 29.4 37.5 51.3 20.8 36.6 36.7 31.7 26.2 29.4 15.4 26.5 39.2 11.9 32.50
+NeCo 40.6 25.7 58.0 20.4 37.1 33.3 37.2 56.2 24.6 37.7 40.8 33.1 23.7 35.1 22.8 23.9 45.4 22.4 34.34
4.3 Ablation Studies
Here, we examine the essential parameters of our method by training NeCo on Pascal VOC12 and ADE20k. We assess its ability to perform linear segmentation and in-context scene understanding using the frozen representations learned with each set of parameters. For in-context scene understanding evaluations, we use $\frac{1}{128}$ fraction of the training data and reduce the spatial dimension to $448^{2}$. The number of training epochs for linear segmentation evaluations is set to 20 epochs. For more ablations, including the effect of training epochs and sorting hyperparameters, refer to Appendix B.
Patch Selection Approach. We demonstrate the effect of selecting patches from the foreground, background, or both in 6(a). Foreground patches are selected using the attention map averaged across heads. Our results indicate that selecting patches from the foreground gives 1% better results compared to the background selection in 3 out of 4 metrics. However, the performance peaks when we select patches from both locations. This improvement can be attributed to the use of scene-centric images for training, where the background often contains meaningful objects that contribute to enhanced performance.
Utilizing a Teacher. We ablate the role of teacher-student architecture in 6(b). As shown, employing a teacher network updated by exponential moving average can significantly improve the performance across all the metrics by 8% to 20%. This is consistent with the previous works (Caron et al., 2021; Grill et al., 2020), which reported a more stable training process when the teacher-student architecture is employed.
Nearest Neighbor Selection Approach. We evaluate the influence of picking nearest neighbors from the same image (intra) or different batch images (inter) and find that selecting patches across images consistently boosts performance by roughly 0.4% to 1% across different metrics. The higher diversity of patches involved in the latter approach likely accounts for this improvement (see Appendix B.7 for full tables).
Training Dataset. 6(d) presents the impact of the training dataset based on the ImageNet-100, Pascal, and COCO datasets. ImageNet-100 comprises relatively simple images with few objects, whereas Pascal and COCO feature more complex scenes with multiple objects. Our method shows consistent improvements when trained on multi-object datasets, achieving a performance increase of 2% to 7% on COCO compared to ImageNet-100. This improvement is due to the greater quantity and diversity of objects per batch in multi-object scenes, which provide stronger learning signals by requiring discrimination against a higher number of objects. Notably, the additional performance boost we observe from finetuning DINOv2R on Pascal—despite it already being trained on this dataset (Oquab et al., 2023)—further underscores the efficacy of our proposed loss function.
Sorting Algorithm. We ablate the effect of changing the sorting algorithm and find that our method maintains strong performance across various approaches, achieving the best results with Bitonic sorting, which slightly outperforms the alternatives on average. Additionally, we investigate the absence of a sorting component, which leads to deteriorated performance. The complete tables for this study, along with an ablation on the sorting steepness parameter, are provided in Appendix B.7, as our method is robust to variations in sorting parameters.
Batch Size. 6(e) examines how batch size affects performance. Smaller batch sizes provide marginal gains, but larger batch sizes show more significant improvements, indicating that performance could be further improved with batch sizes exceeding 64.
Number of Neighbors. As detailed in the method section, we use a differentiable sorting algorithm to compute and sort the distances between each patch and others. The ablation study in 6(c) evaluates selecting the top $K$ distances instead of all. Results show that incorporating more neighbors improves performance, but beyond a threshold (e.g., 32), the effect diminishes, indicating robustness against this hyperparameter.
Table 6: Ablations of the key parameters of our method. We evaluate the models by training a linear layer on top of the frozen representations (Lin.) or using the in-context (IC) evaluation of Balazevic et al. (2023) using the validation images for PascalVOC12 and ADE20k.
@@ 4.3 Ablation Studies
Batch Size Lin. IC Lin. IC
1 76.0 60.0 35.4 20.1
4 76.2 60.2 35.6 20.2
8 76.8 61.1 36.3 20.8
16 77.7 61.2 36.7 21.1
32 78.2 61.4 37.1 21.4
64 78.9 62.0 37.3 21.7
5 Conclusion
In this work, we propose Patch Nearest Neighbor Consistency as a new method for dense post-pretraining of self-supervised backbones. By applying our method to the many backbones including the DINOv2-registers model, we improve upon these models by a large margin for frozen clustering, semantic segmentation, full finetuning, and 3D understanding, setting several new state-of-the-art performances.